The December 31, 2017, adjusted trial balance of Maritime Manufacturing showed the following information $ 740,000 360,300 107,600 58,600 Machinery Accumulated depreciation, machinery Office furniture Accumulated depreciation, office furniture2 Remaining useful life four years; estimated residual $90,000 Remaining useful life five years, estimated residual $13,000 Early in 2018, the company made a decision to stop making the items produced by the machinery and buy the items instead As a result, the remaining useful life was decreased to two years and the residual value was increased to a total of $140,000 At the beginning of 2018, it was determined that the estimated life of the office furniture should be reduced by two years and the residual value decreased by $9,000. The company calculates depreciation using the straight-line method to the nearest month Required Prepare the entries to record depreciation on the machinery and office furniture for the year ended December 31, 2018. (If no entry is required for a transactionlevent, select No journal entry required in the first account field. Round the final answer to the nearest whole dollars.) View transaction list Journal entry worksheet 2 Record the depreciation on the machinery, for the year ended Dec. 31, 2018 Note: Enter debits before credits Date General Journal Debit Credit Dec. 31, 2018 Record entry Clear entry View general journal
Answer #1
Depreciation under straight-line method=(Cost-residual value)/Useful life
Machinery:
Machiery cost to be depreciated for the remaining two years=(Cost-accumulated depreciation)=740000-360300=$ 379700
Residual value=$ 140000
Depreciation for 2018=(379700-140000)/2=$ 119850
Journal entry to record depreciation:
Date General journal Debit Credit
Dec 31,2018 Depreciation expense 119850
Accumulated depreciation-Machinery 119850
Office furniture:
Office furniture cost to be depreciated for the remaining years=(Cost-accumulated depreciation)=107600-58600=$ 49000
Residual value=13000-9000=$ 4000
Remaining useful life=5-2=3 years
Depreciation for 2018=(49000-4000)/3=$ 15000
Journal entry to record depreciation:
Date General journal Debit Credit
Dec 31,2018 Depreciation expense 15000
Accumulated depreciation-Office furniture 15000